# ACE: A general-purpose non-Markovian open quantum systems simulation toolkit based on process tensors

### v1

@@ V Summary
The ACE code implements several methods to calculate PT-MPOs and allows PT-MPOs to be stored in files, manipulated, combined with other PT-MPOs, and read again to efficiently scan simulation results for different time-dependent system Hamiltonians. When the environment is of the form of a Gaussian spin-boson model, one can use the algorithm by Jørgensen and Pollock [29], the divide-and-conquer algorithm of Ref. [40], or periodic PT-MPOs. If the environment is more generally composed of independent modes, the ACE algorithm of Ref. [33] as well as an enhanced version with a tree-like combination scheme [43] is available. Moreover, it is possible to compose open quantum system and use several PT-MPOs to describe environment influences. This makes the code extremely versatile, and it demonstrates the core idea for the development of a universal numerically exact solver for networks of non-Markovian open quantum systems based on PT-MPOs.
@@
A current limitation of the code is the size of the open quantum system network because the propagation of a multi-partite system with several PT-MPOs by the analog of Eq. (7) involves the multiplication of matrices, whose dimensions are the products of the individual system Liouville space dimensions and the inner bonds of the PT-MPOs. Future work will be directed towards tackling the exponential scaling with the number of constituent parts of the quantum networks.
@@ Appendix A Summary of commands and arguments
add_Output matrix Specifies an observable [argument] to be extracted from the reduced system density matrix. Every occurrence of add_Output results in the addition of two columns to the outfile, corresponding to real and imaginary parts of the observable.
@@
Table 2: List of commands that can be specified in configuration files. For more commands for PT-MPO generation see Tab. 3. The second column lists the type and order of arguments of the respective command, which are described in the third column. Multiple arguments are separated by white spaces. Types are either text strings (string; must not contain whitespaces), floating point numbers (float), integers (int), matrix-valued expressions (matrix) as described in Sec. III.2, or boolean values (bool). Floating point numbers can also be specified as 1x1 matrix-valued expressions. Arguments in square brackets are optional.

### v2

@@ V Summary
The ACE code implements several methods to calculate PT-MPOs and allows PT-MPOs to be stored in files, manipulated, combined with other PT-MPOs, and read again to efficiently scan simulation results for different time-dependent system Hamiltonians. When the environment is of the form of a Gaussian spin-boson model, one can use the algorithm by Jørgensen and Pollock [35], the divide-and-conquer algorithm of Ref. [45], or periodic PT-MPOs. If the environment is more generally composed of independent modes, the ACE algorithm of Ref. [39] as well as an enhanced version with a tree-like combination scheme [48] is available. Moreover, it is possible to use several PT-MPOs to describe multiple distinct environment influences. This makes the code extremely versatile, and it demonstrates the core idea for the development of a universal numerically exact solver for networks of non-Markovian open quantum systems based on PT-MPOs.
A current limitation of the code is the size of the open quantum system that can be tackled. For example, the PT-MPO algorithm by Jørgensen and Pollock [35] nominally scales as $\mathcal{O}(D^{8})$ with the system dimension $D$ [^3] Exact SVD decomposition algorithms of a $N\times M$ matrix with $N>M$ generally scales as $\mathcal{O}(NM^{2})$. For a single row of the tensor network in the Jørgensen-Pollock approach [35], $M$ corresponds to the inner bonds dimension $D^{2}$, while $N=MD^{2}$ additionally contain the outer bonds. This yields $\mathcal{O}(D^{8})$ overall. For non-Gaussian environments with general off-diagonal coupling to modes of Hilbert space dimension $D_{M}$, the nominal scaling of single SVDs in ACE is $\mathcal{O}(D^{4}D_{M}^{6})$. However, whenever the system-environment coupling operator has degeneracies, such as for composite systems, the outer bond dimension is drastically reduced as described in Sec. II.4. For Gaussian methods, the automatic detection of degeneracies also reduces in inner bond dimension of the initial tensor network. This way, it was possible to use our ACE code for simulations of open quantum systems with over 30 levels, e.g., 4 levels of a QD coupled to two bosonic modes truncated each at Hilbert space dimension 3 in Ref. [71] and superradiant emission from 5 closely spaced quantum emitters in Ref. [45]. For situations where no such degeneracies can be exploited, an approach along the lines of Ref. [72] may enable the simulation of larger system by systematically approximating the coupling Hamiltonian by a lower-rank operator using Chebyshev interpolation.
It is worth stressing once more that a key feature distinguishing PT-MPOs from other approaches is the former’s independence of the system propagator. On the one hand, this enables fast scans of system parameters with a PT-MPO that has to be calculated only once, the evaluation of arbitrary multi-time correlation functions, and the numerically exact simulation for systems with multiple non-Markovian environments. On the other hand, this feature prohibits PT-MPO methods from utilizing assumptions about the system dynamics, which would limit the complexity of the allowed environment response. This is manifested in the fact that the inner bond dimensions $\chi$ of PT-MPOs play a crucial role resulting in an additional $\mathcal{O}(\chi^{3})$ scaling, e.g., for system propagation, where the value of $\chi$ strongly depends on the concrete environment and is difficult to estimate a priori. Thus, for large spin-boson systems coupled to strongly structured environments (and consequently large bond dimensions) and with a fixed, time-independent system Hamiltonians, methods based on the construction of effective propagators operating only on the reduced system Liouville space like SMatPI [29] likely outperform PT-MPOs. For time-independent Hamiltonians, effective propagators for long-time dynamics can also be extrapolated from short-time propagators (over about $n_{\textrm{mem}}$ time steps) from any non-Markovian open quantum systems approach by post-processing using the Transfer Tensor Method (TTM) [73]. To offer access to the improved performance that can become available once a system Hamiltonian has been fixed, our code also provides implementations of QUAPI and TEMPO as alternatives to PT-MPOs.
@@
Finally, note that the propagation of multi-partite system with several PT-MPOs by the analog of Eq. (7) involves the multiplication of matrices, whose dimensions are the products of the individual system Liouville space dimensions and the inner bonds of the PT-MPOs. Future work will be directed towards tackling the exponential scaling in simulations of quantum networks with respect to the number of constituent parts. The combination of PT-MPOs with many-body techniques like TEBD in Ref. [43] exemplifies promising efforts in this direction.
@@ Appendix A Summary of commands and arguments
set_precision int Changes number of significant digits of floating point numbers written to outfile
@@
add_Output matrix Specifies an observable [argument] to be extracted from the reduced system density matrix. Every occurrence of add_Output adds two columns to the outfile, corresponding to real and imaginary parts of the observable.
@@
threshold_range_factor float The threshold is multiplied by a factor, which is exponentially interpolated from $1/r$ for the first MPO compression sweep to $1$ for the final compression sweep.
@@
Table 2: Commands that can be specified in configuration files. For more commands for PT-MPO generation see Tab. 3. The second column lists the type and order of arguments. A description is provided in the third column. Multiple arguments are separated by white spaces. Types are either text strings (string; must not contain whitespaces), floating point numbers (float), integers (int), matrix-valued expressions (matrix) as described in Sec. III.2, or boolean values (bool). Floating point numbers can also be specified as 1x1 matrix-valued expressions. Arguments in square brackets are optional.